OpenStudy (idealist10):

How to integrate (1+2z)/(1-z^2) dz.

ganeshie8 (ganeshie8):

hint : 1-z^2 = (1+z)(1-z)

OpenStudy (idealist10):

So partial fractions?

OpenStudy (freckles):

You could use partial fractions or you could write \[\frac{1+z+z}{1-z^2}=\frac{1+z+z}{(1-z)(1+z)}=\frac{1+z}{(1-z)(1+z)}+\frac{z}{1-z^2}\] but partial fractions works too